OpenStudy (anonymous):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 3, – 25), and x intercepts at x = -8 and x = 2. (Do not include the negative sign in your answer.) y = x2 + 6x − ___

OpenStudy (anonymous):

x intercepts means y = 0. So, when x = -8 or x = 2, y = 0. Then solve for the unknown constant

OpenStudy (anonymous):

\[y = x^2 + 6x + 9 \rightarrow y = x^2 + 6x + 3^2\]

OpenStudy (anonymous):

or \[y=x^2+6x+9→y=x^2+6x+(-3)^2\]

OpenStudy (anonymous):

Thats wrong

OpenStudy (anonymous):

ugh im so confused :(

OpenStudy (anonymous):

What i meant is substitute x = 2 or x = -8(your choice) into the equation and y = 0. Then solve for the constant

OpenStudy (anonymous):

so would it look like 0=-8^2+6(2)

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Just choose one x

OpenStudy (anonymous):

can you show me?

OpenStudy (anonymous):

0 = (2)^2 + 6(2) - __

OpenStudy (anonymous):

Or 0 = (-8)^2 + 6(-8) - __ you will get the same answer

OpenStudy (anonymous):

so how do i find the missing number?

OpenStudy (anonymous):

0 = 2^2 + 6(2) - __ 0 = 4 + 12 - __ 16 - __ = 0 __ = 16

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

Yes